OpenStudy (anonymous):

14^x+1=36 Round to the nearest ten thousandth

OpenStudy (kirbykirby):

\[\large 14^x+1=36\\ \large 14^x=36-1\\ \large 14^x=35\\ \large \log(14^x)=\log 35\\ \large x \log14 = \log 35\\ \large x = \frac{\log35}{\log 14}\approx 1.3472\]

OpenStudy (kirbykirby):

actually maybe you meant instead \( \large 14^{x+1}=36\) ? \[ \large 14^{x+1}=36\\ \large \log(14^{x+1})=\log 36\\ \large (x+1)\log 14 = \log 36\\ \large x \log14 + \log 14 = \log 36\\ \large x\log 14 = \log 36 - \log 14\\ \large x = \frac{\log 36 - \log 14}{\log 14}\approx 0.3579\]